OpenStudy (ebonyserrano):

A copy machine makes 28 copies per minute. How many copies does it make in 4 minutes and 30 seconds?

jhonyy9 (jhonyy9):

one minute how many second have ?

jhonyy9 (jhonyy9):

ebony this is same with the past exercise

OpenStudy (ebonyserrano):

60

jhonyy9 (jhonyy9):

one minute has 60 seconds so now you know that this machine make 28 copies in 60 seconds 28 copies in 60 sec. x copies in 4 min. 30 sec. 4 min. 30 sec. equal 4*60 +30 = 240 +30 = 270 sec. 28 copies ----- 60 sec. x copies ----- 270 sec. ------------------------ x = 270*28/60 = ? sec.

jhonyy9 (jhonyy9):

sorry x will get in copies not sec. i confused it with the past exercise so this mean that you get with this result exactly the right answer easy

OpenStudy (ebonyserrano):

so 126?

jhonyy9 (jhonyy9):

yes 126 copies

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

do you have questions ?